# Thermal dilepton polarization and dynamics of the QCD plasma in relativistic heavy-ion collisions

### v1

@@ Abstract
We observe an overall sign change and a significant enhancement in $\lambda_{\theta}$ when going from LO to NLO. This better matches the physical expectation that at small invariant mass, the virtual photon behaves like a real photon with two transverse polarization modes. In contrast, the LO calculation does not capture this aspect due to the vanishing phase space for $M\to 0\,$, where the NLO result is actually the “leading” contribution. At large invariant masses, the NLO result is a small correction to the LO one and both reflect the lack of a preferred polarization direction for $M\gg T$. Underlying this dramatic difference between orders in the perturbative calculation, is the participation of (thermal) gluons in the production reactions. The importance of gluons is readily supported from non-perturbative lattice studies, where the current-current correlators do not change qualitatively when going from quenched to $2$+$1$ flavour QCD Brandt et al. (2018); Cè et al. (2022); Ali et al. (2024).
@@
We also explore the dilepton polarization phenomenon induced during the pre-equilibrium stage and the QGP phase. It is found that dilepton polarization can be highly sensitive to the pre-equilibrium evolution; we find that the effects of the pre-equilibrium phase on dilepton polarization become more pronounced when considering the transverse momentum $p_{{}_{\rm T}}$-dependence in the IMR. Therefore, studies of the pre-equilibrium stages in heavy-ion collisions will benefit from those of dilepton polarization.

### v3

@@ Abstract
We observe an overall sign change and a significant enhancement in $\lambda_{\theta}$ when going from LO to NLO. This better matches the physical expectation that at small invariant mass, the virtual photon behaves like a real photon with two transverse polarization modes. In contrast, the LO calculation does not capture this aspect due to the vanishing phase space for $M\to 0\,$, where the NLO result is actually the “leading” contribution. At large invariant masses, the NLO result is a small correction to the LO one and both reflect the lack of a preferred polarization direction for $M\gg T$. Underlying this dramatic difference between orders in the perturbative calculation, is the participation of (thermal) gluons in the production reactions. The importance of gluons is readily supported from nonperturbative lattice studies, where the current-current correlators do not change qualitatively when going from quenched to $2$+$1$ flavor QCD Brandt et al. (2018); Cè et al. (2022); Ali et al. (2024).
@@
We also explore the dilepton polarization phenomenon induced during the preequilibrium stage and the QGP phase. It is found that dilepton polarization can be highly sensitive to the preequilibrium evolution; we find that the effects of the preequilibrium phase on dilepton polarization become more pronounced when considering the transverse momentum $p_{{}_{\rm T}}$-dependence in the IMR. Therefore, studies of the preequilibrium stages in heavy-ion collisions will benefit from those of dilepton polarization.